OpenStudy (nuccioreggie):

A jar contains 0.75 liter of blackberry juice and 0.60 liter of blueberry juice. Patrick poured 0.25 liter of guava juice into the jar. He then drank 0.20 liter of the mixture. Part A: Write an expression to represent the total amount of juice left in the jar. (5 points) Part B: Identify which property is used in each step to simplify the expression. (5 points)

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

hint: total liters into the jar are: 0.75+0.60+0.25=..?

OpenStudy (michele_laino):

what is: 0.75+0.60+0.25=...?

OpenStudy (nuccioreggie):

1.6

OpenStudy (michele_laino):

ok! Now we have subtract 0.2 liters, so we can write: 1.6-0.2=...?

OpenStudy (michele_laino):

what is 1.6-0.2=...?

OpenStudy (nuccioreggie):

Hey i will brb ok 2mins

OpenStudy (nuccioreggie):

Back

OpenStudy (michele_laino):

ok! what is 1.6-0.2=...?

OpenStudy (nuccioreggie):

1.4

OpenStudy (michele_laino):

ok! that's right. That's your answer to the first part of your question

OpenStudy (nuccioreggie):

Ok thanks now we can do the next one

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

at the first step, we have applied the associative property of addition, since we have wrote this: (0.75+0.60)+0.25=1.60

OpenStudy (nuccioreggie):

Ok

OpenStudy (michele_laino):

at the second step, we have computed a subtraction, so we have applied the rule in order to perform a subtraction

OpenStudy (nuccioreggie):

Ok so and is thAt one of the things i write down also??

OpenStudy (michele_laino):

yes!